40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3162     Entry time: Tue Jul 6 17:38:27 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Tip Tilt Progress! 

[Jenne, Kyung Ha]

We successfully suspended the 4 eddy current dampers for the first Tip Tilt.  We had some lessons learned, including how to carefully get an allen wrench in between the dampers to do up some of the screws, and how to be careful not to bend the wire while tightening the screws.  More tomorrow...

ELOG V3.1.3-